OpenStudy (anonymous):

The rocket-driven sled Sonic Wind No. 2, used for investigating the physiological effects of large accelerations, runs on a straight, level track that is 1080 m long. Starting from rest, it can reach a speed of 1560 km/h in 1.70 s. a) Compute the acceleration in m/s^2. b) Compute the acceleration in g’s. c) What is the distance covered in 1.70s?

OpenStudy (anonymous):

a)v=u(starting speed which in this case =0)+at for converting the km/h to m/s multiply by 1000/3600 b)1 meter / (second^2) = 0.101971621 g c)s=ut(again 0) +at^2/2

OpenStudy (anonymous):

for c distance = 0 + (26(1.7)6^2)/2 right?

OpenStudy (anonymous):

d= 0+(26/2)(1.7^2)

OpenStudy (anonymous):

yep, if the acc is 26m/s^2, i haven't checked, yet your logic is right.

OpenStudy (anonymous):

do you know what a is?

OpenStudy (anonymous):

acceleration

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cos i dont think 26 is the accerelation

OpenStudy (anonymous):

a)Okay, first you have that 1560 km/h = 433.333333 meters / second. Then using v=at you find that the acceleration is 433.333333/1.7 = 254.901961m/s^2 b)1 meter / (second^2) = 0.101971621 g c)s=254.901/2*(1.7)^2 m

OpenStudy (anonymous):

thank you